OpenStudy (anonymous):

Peanuts sell for $3 per pound at a local store. Chocolate candies sell for $6 per pound at the same store. How many pounds of chocolate candies must be purchased to have a 5–pound mixture that costs $4.20 per pound? Answer 3 2 4 1 Peanuts sell for $3 per pound at a local store. Chocolate candies sell for $6 per pound at the same store. How many pounds of chocolate candies must be purchased to have a 5–pound mixture that costs $4.20 per pound? Answer 3 2 4 1 @Mathematics

OpenStudy (anonymous):

answer is 2 i didn't read the per pound part so you have to find 4.20 x 5 so it'd be 3x+6y=5*4.20 x+y=5

OpenStudy (anonymous):

solving this equation gets 3 peanut pounds and 2 chocolate pounds

hero (hero):

Seems weird to me

hero (hero):

The chocolate already costs $6 per pound. Why would you buy two of them?

OpenStudy (anonymous):

because you are buying 5 pounds of mixture.. so you are essentially buying 4.20 * 5 of the mixture

hero (hero):

That's not a good enough explanation

OpenStudy (anonymous):

which is 21 dollars... i was getting negative answers the first time lol since i did if it all cost 4.20 lol

OpenStudy (anonymous):

3(3)+2(6)=21? is that good enough explanation?